Soso brought two cats, Gretchen and Tk, from a pet store in 2015. Gretchen's weight was 3 killograms, and Tk 5 kilograms. In 2018, after a period of growth, Gretchen weight 6 kilograms and Tk 8 kilograms.

1. Which cat do you think showed the largest growth, and why?

2. Tawanda, Soso's friend,says that neither of the cats had growth more than the other because both cats had gained precisely 1 kilogram in weight.

3. How did Tawanda determine this methematically? Write down the computations the Tawanda might have made.


1
Expert's answer
2021-06-28T16:38:50-0400

growth rate = present-past/past


Gretchen's growth rate = "\dfrac{6-3}{3} = 1"


TK's growth rate = "\dfrac{8-5}{5} = 0.6"


1. "\therefore" Gretchen showed the largest growth

2. Tawanda isn't analytically correct

3. Tawanda determined his/her inference by subtracting the cats old weight from their new/present weight.